### v1

@@ Abstract
keywords:
[…]
1 Introduction
Due to global warming, forest fires are becoming increasingly frequent, causing severe environmental damage and posing threats to human life and property safety. To mitigate the risk of forest fires, forest fire prevention efforts must not only focus on proactive prevention and management but also emphasize the rapid detection, early warning, and timely response to fires. In forest scenarios, smoke is a critical early indicator of fire. Smoke detection based on deep learning has emerged as a significant approach for early forest fire prevention. However, deep learning-based smoke detection models require a large and diverse dataset of smoke images. Currently, publicly available datasets for forest fire smoke are limited in quantity and quality, making it challenging to train highly accurate and generalizable smoke detection models.
Currently, there are two primary methods for acquiring forest fire smoke image data. The traditional approach involves collecting smoke data generated from real combustion, such as image records of actual forest fires or photographs of smoke produced during controlled ignition experiments. However, obtaining early-stage smoke images from real forest fires is extremely challenging. For controlled ignition experiments, if conducted in laboratory settings, the resulting datasets are limited to laboratory backgrounds, which lack diversity and lead to poorly generalizable models. On the other hand, conducting experiments in outdoor settings, such as forests, woodlands, or nature reserves, is often impractical due to high risks and costs, making it difficult to obtain large-scale fire smoke data. Due to the aforementioned reasons, the data currently used for smoke detection generally suffer from low quality, as shown in Figure 1. Therefore, relying solely on collecting smoke images from real combustion cannot fully meet the practical needs of forest fire detection. With the widespread application of digital imaging technology and deep learning, using image processing techniques or deep learning models to synthesize smoke data has become a crucial method for data acquisition. Image processing techniques primarily involve manipulating smoke in existing images through simulation or modeling software to generate synthetic images. Deep learning, on the other hand, focuses on learning features from existing images to generate new images.
@@ 1 Introduction
In the field of smoke and fire detection, the use of synthetic data currently relies primarily on image processing techniques and generative adversarial networks (GANs), with a smaller portion utilizing diffusion models for data synthesis. Methods based on image processing (Labati et al. (2013), Xu et al. (2017), Yuan et al. (2019), Mao et al. (2021), Wang et al. (2022)) mainly involve rendering and simulating realistic smoke and fire data using various synthesis software and modeling techniques. These approaches have been applied to early-stage smoke and fire image synthesis and integrated into detection models. Recently, some studies have employed Unreal Engine 5 (Wang et al. (2025), Wang et al. (2024a)) to generate forest wildfire datasets, which are subsequently used for training detection models. The use of generative adversarial networks (GANs) (Namozov & Im Cho (2018), Huo et al. (2024)) for synthesizing smoke and fire images has become a common method in recent years. These methods typically involve designing generators and discriminators for adversarial training to produce smoke and fire data similar to the dataset or to generate controllable smoke images by providing mask-based conditional inputs. Approaches based on diffusion models (Wang et al. (2024b), Zheng et al. (2024a), Zheng et al. (2024b)) for generating smoke and fire data are still relatively few. Their design principles are similar to those of GANs, involving fine-tuning pre-trained diffusion models to generate smoke data with similar distributions or enabling controllable synthesis through editing techniques.
@@
Although traditional image processing techniques can generate smoke images, the fusion of the background and smoke regions is relatively simplistic, resulting in limited realism in the synthesized images. Consequently, datasets composed of such synthetic data offer only limited improvements in the accuracy and generalization of detection models. Generative Adversarial Networks (GANs), introduced by GoodFellow et al. (Goodfellow et al. (2014)) in 2014, have become a pivotal technology in the field of image generation. However, GANs suffer from issues such as training instability, gradient vanishing, and mode collapse. Additionally, in the synthesis of smoke, GANs tend to produce noticeable boundaries, leading to inconsistencies in the context of the synthesized images. Diffusion models, the state-of-the-art in image generation, were first proposed by Ho et al. (Ho et al. (2020)) in 2020. These models generate images with superior diversity and realism and allow for guided and controlled outputs through various conditions, surpassing previous image synthesis methods. The use of diffusion models for dataset augmentation has already found applications in general image editing, medical image processing, and artistic creation. Despite their impact across multiple domains, the application of diffusion models in the field of forest fire smoke remains limited. Moreover, due to the suboptimal performance of earlier diffusion-based text-to-image models in generating smoke data, which often produce erroneous and unrealistic results, and the inability of image inpainting models to achieve sufficient realism and contextual consistency in controllable smoke generation, pre-trained diffusion models cannot be directly utilized for generating forest smoke images. Furthermore, given the diversity and uncertainty inherent in diffusion model outputs, it is necessary to filter out high-quality smoke images. Therefore, there is a need to design a method specifically tailored for forest smoke generation to improve the synthesis efficiency and overall quality of smoke datasets.
@@
To address the aforementioned issues, we proposed a novel generation method tailored for smoke data, designed to generate forest smoke images with specified backgrounds and smoke contours. Additionally, we implemented a filtering process for the synthesized smoke data to enhance the quality of the dataset and improve the performance of smoke detection models. Specifically, the main contributions of this paper are as follows:
@@
(1) We proposed an automated pipeline for generating synthetic smoke and fire datasets. By leveraging the pre-trained segmentation model SAM (Kirillov et al. (2023)), we automatically generated masks for the smoke regions based on previously annotated data used for training detection models. Additionally, we employed the multimodal model Blip2 (Li et al. (2023)) to generate captions for the smoke images. After organizing and pairing these components, we obtained a dataset for training smoke image inpainting models.
@@
(3) We proposed a data filtering mechanism tailored to the characteristics of smoke. By fine-tuning an advanced multimodal large language model based on smoke properties, we filtered out high-quality smoke datasets. These datasets are then automatically annotated using image masks and subsequently used to train detection models.
The rest of this paper is organized as follows. In Section 2, we provide an overview of related research relevant to this study. In Section 3, we introduce some preliminary knowledge associated with our proposed method. Section 4 details the methodology we propose. In Section 5, we describe the experimental procedures and present the results. Finally, the concluding remarks are provided in the last section 6.
2 Related work
@@
Fire smoke synthesis. Labati et al. (2013) synthesized wildfire smoke images using a cellular model virtual environment and image processing techniques, which were then used to train and test outdoor smoke detection algorithms. Xu et al. (2017) employed Blender to simulate and render smoke images and utilized domain adaptation techniques to align the feature distributions of synthetic and real smoke images, thereby improving detection performance. Zhang et al. (2018a) generated forest smoke images by inserting real or simulated smoke into forest backgrounds and used these synthetic images to train a Faster R-CNN model for forest fire smoke detection. Namozov & Im Cho (2018) addressed the overfitting issue caused by training networks on limited datasets by using traditional data augmentation techniques and generative adversarial networks (GANs) to increase the number of available training smoke images. Wang et al. (Wang et al. (2024a), Wang et al. (2025)) enhanced the dataset by synthesizing multi-angle forest wildfire data using Unreal Engine 5 and trained their proposed detection models on this dataset. Yuan et al. (2019) proposed a method for synthesizing smoke to tackle the challenge of annotating ambiguous smoke images. Mao et al. (2021) simulated a large variety of richly shaped smoke using 3D modeling software and rendered virtual smoke in numerous virtual wilderness background images with diverse environmental settings, constructing a synthetic dataset to supplement smoke datasets. Wang et al. (2022) improved a YOLOv5-based detection model by collecting a large number of real and synthetic smoke datasets. Huo et al. (2024) enhanced generative adversarial networks with a multi-scale fusion module, using image masks as conditions for controllable smoke image generation. Wang et al. (2024b) proposed a diffusion model-based framework for generating wildfire image data using diffusion models and real flame image masks. Zheng et al. (2024a) introduced a dataset quality enhancement framework based on diffusion models (DDPM) to improve the quality of low-quality fire alarm datasets. Zheng et al. (2024b) utilized pre-trained diffusion models to generate fire images and applied them to create high-quality fire segmentation datasets.
@@
Our proposed method is implemented based on the pre-trained Stable Diffusion inpainting model. Stable Diffusion, an improvement on the Latent Diffusion Model (LDM) introduced by Rombach et al. (2022), consists of three components: a Variational Autoencoder (VAE) (Kingma (2013)), a denoising U-Net (Ronneberger et al. (2015)), and a CLIP (Radford et al. (2021)) text encoder. Although the pre-trained Stable Diffusion inpainting model exhibits powerful image inpainting capabilities, its performance in various derived inpainting tasks is not outstanding. Therefore, research focusing on different forms of image inpainting tasks has become a hotspot in recent years. Currently, methods for image inpainting tasks are primarily divided into two categories: those based on sampling strategy modifications and those based on dedicated inpainting models.
Strategy-based modifications. Strategy-based modifications primarily involve manipulation during the diffusion model’s inference process to achieve inpainting or editing goals. Meng et al. (2021) proposed leveraging the diffusion model’s generative prior and synthesizing realistic images through iterative denoising. Hertz et al. (2022) achieved training-free image editing by adjusting the cross-attention maps between the text embedding space and the image space. Couairon et al. (2022) automatically generated masks based on text queries to highlight regions requiring editing and preserved content through latent inference. Kim et al. (2022) utilized the inversion and generation capabilities of diffusion models to achieve zero-shot image manipulation. Park et al. (2024) introduced object contours as constraints, enhancing control over inpainting tasks. Mokady et al. (2023) proposed an accurate image inversion technique combined with a single pivot noise vector to improve text-guided real image inpainting. Xie et al. (2023a) achieved spatial control over objects and context by adding spatial constraints during the denoising steps. Avrahami et al. (2022) were the first to achieve general natural image inpainting based on natural language descriptions and ROI masks. To optimize computational efficiency, they later introduced the Blended Latent Diffusion(Avrahami et al. (2023)), operating in a low-dimensional latent space to accelerate inference and reduce the need for CLIP gradient calculations.
@@
End-to-End inpainting model. End-to-end methods typically require modifying the model or fine-tuning the original model with control conditions to learn inpainting and various editing operations in an end-to-end manner. Xie et al. (2023b) combined text and shape guidance to provide flexible control over object inpainting. Zou et al. (2024) proposed a method that eliminates the need for image mask inputs, leveraging the cross-modal attention capabilities of pre-trained diffusion models for instant inpainting. Zhuang et al. (2025) improved the quality of multi-task image inpainting through task prompts and customized fine-tuning strategies. Ju et al. (2024) adopted a plug-and-play design, separating occluded image features from noise to reduce the learning burden and facilitate information fusion. Nair (2024) proposed feature-aligned diffusion, enhancing generation quality by reducing the discrepancy between the image encoder output and the diffusion model features. Brooks et al. (2023) introduced the first end-to-end image editing model based on human instructions, using large language models to generate editing instructions and datasets for training. Qi et al. (2024) incorporated the Q-former (Li et al. (2023)) structure to train the model for style transfer capabilities. Wei et al. (2024) integrated multiple editing task models, generated a large number of editing pairs, and trained them on the DiT architecture, achieving unified editing functionality.
3 Preliminaries
@@
In this section, we will introduce the fundamentals of Latent Diffusion Model in Section 3.1. Since end-to-end methods generally achieve better performance compared to strategy-based approaches, we also adopt an end-to-end method for our structural design in this work. Therefore, in Section 3.2, we present some basic methods currently used for end-to-end image inpainting.
3.1 Latent Diffusion Model
@@
The LDM architecture primarily consists of three components: a Variational Autoencoder (VAE) (Kingma (2013)), which compresses images into the latent space and decodes the denoised latent representations back into images. The VAE significantly enhances the generative performance of diffusion-based methods, enabling the model to produce high-resolution images. A denoising U-Net (Ronneberger et al. (2015)), which iteratively refines randomly sampled Gaussian noise through multiple U-Net layers to obtain clean latent representations. A CLIP (Radford et al. (2021)) text encoder, which generates text embeddings that guide the U-Net denoising process via cross-attention mechanisms.
3.2 Previous End-to-End Inpainting Models
@@
4 Methodology
4.1 Data Acquisition
@@
Specifically, to preserve the text-guided capabilities of the pre-trained image inpainting model, we employed the pre-trained image-text multimodal model BLIP-2 (Li et al. (2023)) to generate captions for smoke images. BLIP-2 (Bootstrapping Language-Image Pre-training) is an advanced model based on multimodal learning, capable of understanding complex image content and generating descriptive text or answering questions about images. The captions output by the multimodal model included information about the smoke and background. However, since the target detection dataset contained detailed information such as location and time, which could affect the results of our smoke image generation, we limited the maximum number of tokens generated by the multimodal model to 20 to prevent overly detailed information from being output. Given that the smoke detection dataset included corresponding annotation files, we used the pre-trained segmentation model SAM (Kirillov et al. (2023)) to segment the smoke. SAM is a general-purpose large-scale image segmentation model that can accept inputs such as positive and negative sample points, bounding boxes, and full segmentation. We employed the bounding box information obtained from previous annotation files as input to SAM, thereby obtained the mask for each image. Ultimately, we obtained over 60,000 smoke image-text pairs and corresponding mask images. These were used to train our subsequent smoke generation model. Figure 2 shows a portion of the dataset.
4.2 MFGDiffusion
End-to-End image inpainting methods are typically trained based on pre-trained image inpainting models. In our proposed network architecture, we also leveraged an advanced pre-trained image inpainting model, adopting SD-2-Inpainting (Rombach et al. (2022)) as the foundational model. Since the original image inpainting model cannot fully extract the feature information from the mask and masked image, we introduced a pre-trained ResNet50 (He et al. (2016)) model as our feature extractor. The features output by the feature extractor are aligned with the dimensions of the intermediate representations of the U-Net through a projection layer and are fused with the intermediate representations via cross-attention layers, thereby guiding the denoising process of the U-Net. The architecture of our model is illustrated in Figure 3.

### v2

@@ Abstract
I Introduction
Due to global warming, forest fires are becoming increasingly frequent, causing severe environmental damage and posing threats to human life and property safety. To mitigate the risk of forest fires, forest fire prevention efforts must not only focus on proactive prevention and management but also emphasize the rapid detection, early warning, and timely response to fires. In forest scenarios, smoke is a critical early indicator of fire. Smoke detection based on deep learning [1], [2], [3], [4], [5] has emerged as a significant approach for early forest fire prevention. However, deep learning-based smoke detection models require a large and diverse dataset of smoke images. Currently, publicly available datasets for forest fire smoke are limited in quantity and quality, making it challenging to train highly accurate and generalizable smoke detection models.
Currently, there are two primary methods for acquiring forest fire smoke image data. The traditional approach involves collecting smoke data generated from real combustion, such as image records of actual forest fires or photographs of smoke produced during controlled ignition experiments. However, obtaining early-stage smoke images from real forest fires is extremely challenging. Controlled ignition in laboratories yields background-limited datasets, reducing diversity and generalizability. On the other hand, conducting experiments in outdoor settings, such as forests, woodlands, or nature reserves, is often impractical due to high risks and costs, making it difficult to obtain large-scale fire smoke data,and the scarcity is further exacerbated by the unpredictability of wildfire occurrences, the difficulty of capturing smoke at its initial stages, and the restrictions imposed by environmental protection and safety regulations. Due to the aforementioned reasons, the data currently used for smoke detection generally suffer from low quality, as shown in Figure 1. Therefore, relying solely on collecting smoke images from real combustion cannot fully meet the practical needs of forest fire detection. With the widespread application of digital imaging technology and deep learning, using image processing techniques or deep learning models to synthesize smoke data has become a crucial method for data acquisition. Image processing techniques primarily involve manipulating smoke in existing images through simulation or modeling software to generate synthetic images. Deep learning, on the other hand, focuses on learning features from existing images to generate new images.
@@ I Introduction
In the field of smoke and fire detection, the use of synthetic data currently relies primarily on image processing techniques and generative adversarial networks (GANs), with a smaller portion utilizing diffusion models for data synthesis. Methods based on image processing [6], [7], [8], [9], [10] mainly involve rendering and simulating realistic smoke and fire data using various synthesis software and modeling techniques. These approaches have been applied to early-stage smoke and fire image synthesis and integrated into detection models. Recently, some studies have employed Unreal Engine 5 [11], [12] to generate forest wildfire datasets, which are subsequently used for training detection models. The use of generative adversarial networks (GANs) [13], [14] for synthesizing smoke and fire images has become a common method in recent years. These methods typically involve designing generators and discriminators for adversarial training to produce smoke and fire data similar to the dataset or to generate controllable smoke images by providing mask-based conditional inputs. Approaches based on diffusion models [15], [16], [17] for generating smoke and fire data are still relatively few. Their design principles are similar to those of GANs, involving fine-tuning pre-trained diffusion models to generate smoke data with similar distributions or enabling controllable synthesis through editing techniques.
@@
Although traditional image processing techniques can generate smoke images, the fusion of the background and smoke regions is relatively simplistic, resulting in limited realism in the synthesized images. Consequently, datasets composed of such synthetic data offer only limited improvements in the accuracy and generalization of detection models. Generative Adversarial Networks (GANs), introduced by GoodFellow et al. [18] in 2014, have become a pivotal technology in the field of image generation. However, GANs suffer from issues such as training instability, gradient vanishing, and mode collapse. Additionally, in the synthesis of smoke, GANs tend to produce noticeable boundaries, leading to inconsistencies in the context of the synthesized images. These challenges arise from the sensitivity of adversarial training dynamics and from the difficulty of modeling semi-transparent, fine-grained structures such as smoke while preserving global contextual consistency.Diffusion models, the state-of-the-art in image generation, were first proposed by Ho et al. [19] in 2020. These models generate images with superior diversity and realism and allow for guided and controlled outputs through various conditions, surpassing previous image synthesis methods. The use of diffusion models for dataset augmentation has already found applications in general image editing, medical image processing, and artistic creation. Despite their impact across multiple domains, the application of diffusion models in the field of forest fire smoke remains limited. Moreover, due to the suboptimal performance of earlier diffusion-based text-to-image models in generating smoke data, which often produce erroneous and unrealistic results, and the inability of image inpainting models to achieve sufficient realism and contextual consistency in controllable smoke generation, pre-trained diffusion models cannot be directly utilized for generating forest smoke images. Furthermore, given the diversity and uncertainty inherent in diffusion model outputs, it is necessary to filter out high-quality smoke images. Therefore, there is a need to design a method specifically tailored for forest smoke generation to improve the synthesis efficiency and overall quality of smoke datasets.
@@
To address the aforementioned issues, we proposed a novel generation method tailored for smoke data, designed to generate forest smoke images with specified backgrounds and smoke contours. Additionally, we implemented a filtering process for the synthesized smoke data to enhance the quality of the dataset and improve the performance of smoke detection models.Compared with recent diffusion-based approaches, our method introduces a dedicated architecture and optimization strategy specifically designed for forest smoke generation, addressing the challenges of morphological diversity, background–foreground fusion, and controllable synthesis that previous methods have not fully resolved. Specifically, the main contributions of this paper are as follows:
@@
(1) We proposed an automated pipeline for generating synthetic smoke and fire datasets. By leveraging the pre-trained segmentation model SAM [20], we automatically generated masks for the smoke regions based on previously annotated data used for training detection models. Additionally, we employed the multimodal model Blip2 [21] to generate captions for the smoke images. After organizing and pairing these components, we obtained a dataset for training smoke image inpainting models.
@@
(3) We proposed a data filtering mechanism tailored to the characteristics of smoke. By fine-tuning an advanced multimodal large language model based on smoke properties, we acquired high-quality smoke datasets. These datasets are then automatically annotated using image masks and subsequently used to train detection models.
The rest of this paper is organized as follows. In Section II, we provide an overview of related research relevant to this study. In Section III, we introduce some preliminary knowledge associated with our proposed method. Section IV details the methodology we propose. In Section V, we describe the experimental procedures and present the results. Finally, the concluding remarks are provided in the last section VI.
II Related Work
@@
Fire smoke synthesis. [6] synthesized wildfire smoke images using a cellular model virtual environment and image processing techniques, which were then used to train and test outdoor smoke detection algorithms. [7] employed Blender to simulate and render smoke images and utilized domain adaptation techniques to align the feature distributions of synthetic and real smoke images, thereby improving detection performance. [22] generated forest smoke images by inserting real or simulated smoke into forest backgrounds and used these synthetic images to train a Faster R-CNN model for forest fire smoke detection. [13] addressed the overfitting issue caused by training networks on limited datasets by using traditional data augmentation techniques and generative adversarial networks (GANs) to increase the number of available training smoke images. Wang et al. ([12], [11]) enhanced the dataset by synthesizing multi-angle forest wildfire data using Unreal Engine 5 and trained their proposed detection models on this dataset. [8] proposed a method for synthesizing smoke to tackle the challenge of annotating ambiguous smoke images. [9] simulated a large variety of richly shaped smoke using 3D modeling software and rendered virtual smoke in numerous virtual wilderness background images with diverse environmental settings, constructing a synthetic dataset to supplement smoke datasets. [10] improved a YOLOv5-based detection model by collecting a large number of real and synthetic smoke datasets. [14] enhanced generative adversarial networks with a multi-scale fusion module, using image masks as conditions for controllable smoke image generation. [15] proposed a diffusion model-based framework for generating wildfire image data using diffusion models and real flame image masks. [16] introduced a dataset quality enhancement framework based on diffusion models (DDPM) to improve the quality of low-quality fire alarm datasets. [17] utilized pre-trained diffusion models to generate fire images and applied them to create high-quality fire segmentation datasets.
@@
Our proposed method is implemented based on the pre-trained Stable Diffusion inpainting model. Stable Diffusion, an improvement on the Latent Diffusion Model (LDM) introduced by [23], consists of three components: a Variational Autoencoder (VAE) [24], a denoising U-Net [25], and a CLIP [26] text encoder. Although the pre-trained Stable Diffusion inpainting model exhibits powerful image inpainting capabilities, its performance in various derived inpainting tasks is not outstanding. Therefore, research focusing on different forms of image inpainting tasks has become a hotspot in recent years. Currently, methods for image inpainting tasks are primarily divided into two categories: those based on sampling strategy modifications and those based on dedicated inpainting models. Since End-to-End inpainting models produce more natural results in image inpainting tasks, we implement our approach based on this paradigm.
@@
End-to-End inpainting model. End-to-end methods typically require modifying the model or fine-tuning the original model with control conditions to learn inpainting and various editing operations in an end-to-end manner. [27] combined text and shape guidance to provide flexible control over object inpainting. [28] proposed a method that eliminates the need for image mask inputs, leveraging the cross-modal attention capabilities of pre-trained diffusion models for instant inpainting. [29] improved the quality of multi-task image inpainting through task prompts and customized fine-tuning strategies. [30] adopted a plug-and-play design, separating occluded image features from noise to reduce the learning burden and facilitate information fusion. [31] proposed feature-aligned diffusion, enhancing generation quality by reducing the discrepancy between the image encoder output and the diffusion model features. [32] introduced the first end-to-end image editing model based on human instructions, using large language models to generate editing instructions and datasets for training. [33] incorporated the Q-former [21] structure to train the model for style transfer capabilities. [34] integrated multiple editing task models, generated a large number of editing pairs, and trained them on the DiT architecture, achieving unified editing functionality.
III Preliminaries
@@
In this section, we will introduce the fundamentals of Latent Diffusion Model in Section III-A. Since end-to-end methods generally achieve better performance compared to strategy-based approaches, we also adopt an end-to-end method for our structural design in this work. Therefore, in Section III-B, we present some basic methods currently used for end-to-end image inpainting.
III-A Latent Diffusion Model
@@
The LDM architecture primarily consists of three components: a Variational Autoencoder (VAE) [24], which compresses images into the latent space and decodes the denoised latent representations back into images. The VAE significantly enhances the generative performance of diffusion-based methods, enabling the model to produce high-resolution images. A denoising U-Net [25], which iteratively refines randomly sampled Gaussian noise through multiple U-Net layers to obtain clean latent representations. A CLIP [26] text encoder, which generates text embeddings that guide the U-Net denoising process via cross-attention mechanisms.
III-B Previous End-to-End Inpainting Models
@@
Figure 3: The network architecture of MFGDiffusion. Mask features ($F_{m}$) and masked image features ($F_{M}$) are extracted, projected, and fused with U-Net intermediate features through cross-attention (CACA) and joint cross-attention (JCAJCA). The mask mainly provides positional and morphological cues, while the masked image supplies detailed background information. Together with text embeddings, these features guide the denoising process to generate realistic smoke images.
IV Methodology
IV-A Data Acquisition
@@
Specifically, to preserve the text-guided capabilities of the pre-trained image inpainting model, we employed the pre-trained image-text multimodal model BLIP-2 [21] to generate captions for smoke images. BLIP-2 (Bootstrapping Language-Image Pre-training) is an advanced model based on multimodal learning, capable of understanding complex image content and generating descriptive text or answering questions about images. The captions output by the multimodal model included information about the smoke and background. However, since the target detection dataset contained detailed information such as location and time, which could affect the results of our smoke image generation, we limited the maximum number of tokens generated by the multimodal model to 20 to prevent overly detailed information from being output. Given that the smoke detection dataset included corresponding annotation files, we used the pre-trained segmentation model SAM [20] to segment the smoke. SAM is a general-purpose large-scale image segmentation model that can accept inputs such as positive and negative sample points, bounding boxes, and full segmentation. We employed the bounding box information obtained from previous annotation files as input to SAM, thereby obtained the mask for each image. Ultimately, we obtained over 60,000 smoke image-text pairs and corresponding mask images. These were used to train our subsequent smoke generation model. Figure 2 shows a portion of the dataset.
IV-B MFGDiffusion
End-to-End image inpainting methods are typically trained based on pre-trained image inpainting models. In our proposed network architecture, we also leveraged an advanced pre-trained image inpainting model. Because the original inpainting approach struggles to fully extract feature information from the mask and masked image, we adopted SD-2-Inpainting [23] as the foundational model. To further enhance feature extraction, we introduced a pre-trained ResNet50 [35] model as our feature extractor. The features output by the feature extractor are aligned with the dimensions of the intermediate representations of the U-Net through a projection layer and are fused with the intermediate representations via cross-attention layers, thereby guiding the denoising process of the U-Net. The architecture of our model is illustrated in Figure 3.